A submerged scuba driver hear the sound of a boat on directly above her on the surface of the lake. at the same time, a friend on dry land 22.0m from the board also hears the horn. The horn is 1.2 m above the surface of the water. what is the distance ( labeled by "?") from the horn to the driver? both air and water are at 20 degree C.
1
Expert's answer
2020-06-08T10:17:15-0400

The following fact will help us to solve the problem: the time that the dry friend measured is equal to the time the diver measured. Say that the diver is h meters below the surface of the water. So, first, express the time for the dry friend:


"t=\frac{D}{v_a},"

where "v_a" is the speed of sound in air.

For the submerged friend, this time will be expressed another way:


"t=\frac{H}{v_a}+\frac{h}{v_w},"

where "v_w" is the speed of sound in water.

Express the depth:


"h=v_w\bigg(t-\frac{H}{v_a}\bigg)."

Substitute the time expressed in the first equation:


"h=\frac{v_w}{v_a}(D-H)=\frac{1482}{343}(22-1.2)=89.9\text{ m}."